OpenStudy (anonymous):

A circle has a diameter of 16 meters. What is the area of the circle?

OpenStudy (anonymous):

64pi

OpenStudy (hoblos):

area = pi*r^2 r=d/2=8 then area= 64pi

OpenStudy (rulnick):

Don't forget units. Answer is 16 pi m^2, or about 50.2655 m^2.